OpenStudy (anonymous):

Rewrite the equation in Ax + By = C form. Use integers for A, B, and C. y+3 = -6(x-1)

OpenStudy (astrophysics):

Try working on the right side of the equal sign, any idea what you can do?

OpenStudy (anonymous):

-6x-6 ?

OpenStudy (astrophysics):

Exactly! Now collect like terms

OpenStudy (anonymous):

subtract -6?

OpenStudy (astrophysics):

Write the whole equation out, as I'm not sure what you mean

OpenStudy (anonymous):

y+3 = -6x-6

OpenStudy (anonymous):

would u subract 3

OpenStudy (astrophysics):

Yes good, keep going

OpenStudy (anonymous):

y = -6x-6 then you subtract -6 and 3?

OpenStudy (astrophysics):

You mean, y+3=-6x-6 then y=-6x-6-3

OpenStudy (astrophysics):

So we have y =-6x-9 right?

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

Alright what's next

OpenStudy (anonymous):

not sure

OpenStudy (astrophysics):

We want it in Ax+By=C form, a,b, and c are coefficients

OpenStudy (astrophysics):

c is just a constant, in our case -9

OpenStudy (astrophysics):

So we just need to move the -6x to the left side right?

OpenStudy (anonymous):

yes

OpenStudy (astrophysics):

Ok so can you do that